# GaLore: Memory-Efficient LLM Training by Gradient Low-Rank Projection

### v1

@@ Abstract
Training Large Language Models (LLMs) presents significant memory challenges, predominantly due to the growing size of weights and optimizer states. Common memory-reduction approaches, such as low-rank adaptation (LoRA), add a trainable low-rank matrix to the frozen pre-trained weight in each layer, reducing trainable parameters and optimizer states. However, such approaches typically underperform training with full-rank weights in both pre-training and fine-tuning stages since they limit the parameter search to a low-rank subspace and alter the training dynamics, and further, may require full-rank warm start. In this work, we propose Gradient Low-Rank Projection (GaLore), a training strategy that allows full-parameter learning but is more memory-efficient than common low-rank adaptation methods such as LoRA. Our approach reduces memory usage by up to 65.5% in optimizer states while maintaining both efficiency and performance for pre-training on LLaMA 1B and 7B architectures with C4 dataset with up to 19.7B tokens, and on fine-tuning RoBERTa on GLUE tasks. Our 8-bit GaLore further reduces optimizer memory by up to 82.5% and total training memory by 63.3%, compared to a BF16 baseline. Notably, we demonstrate, for the first time, the feasibility of pre-training a 7B model on consumer GPUs with 24GB memory (e.g., NVIDIA RTX 4090) without model parallel, checkpointing, or offloading strategies.
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive performance across multiple disciplines, including conversational AI and language translation. However, pre-training and fine-tuning LLMs require not only a huge amount of computation but is also memory intensive. The memory requirements include not only billions of trainable parameters, but also their gradients and optimizer states (e.g., gradient momentum and variance in Adam) that can be larger than parameter storage themselves  (Raffel et al., 2023; Touvron et al., 2023; Chowdhery et al., 2022). For example, pre-training a LLaMA 7B model from scratch with a single batch size requires at least 58 GB memory (14GB for trainable parameters, 42GB for Adam optimizer states and weight gradients, and 2GB for activations [^1] The calculation is based on LLaMA architecture, BF16 numerical format, and maximum sequence length of 2048.). This makes the training not feasible on consumer-level GPUs such as NVIDIA RTX 4090 with 24GB memory.
Figure 1: Memory consumption of pre-training a LLaMA 7B model with a token batch size of 256 on a single device, without activation checkpointing and memory offloading. Details refer to Section 5.5.
@@
Parameter-efficient fine-tuning (PEFT) techniques allow for the efficient adaptation of pre-trained language models (PLMs) to different downstream applications without the need to fine-tune all of the model’s parameters (Ding et al., 2022). Among them, the popular Low-Rank Adaptation (LoRA Hu et al. (2021)) reparameterizes weight matrix $W\in\mathbb{R}^{m\times n}$ into $W=W_{0}+BA$, where $W_{0}$ is a frozen full-rank matrix and $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$ are additive low-rank adaptors to be learned. Since the rank $r\ll\min(m,n)$, $A$ and $B$ contain fewer number of trainable parameters and thus smaller optimizer states. LoRA has been used extensively to reduce memory usage for fine-tuning in which $W_{0}$ is the frozen pre-trained weight. Its variant ReLoRA is also used in pre-training, by periodically updating $W_{0}$ using previously learned low-rank adaptors (Lialin et al., 2023).
However, many recent works demonstrate the limitation of such a low-rank reparameterization. For fine-tuning, LoRA is not shown to reach a comparable performance as full-rank fine-tuning (Xia et al., 2024). For pre-training from scratch, it is shown to require a full-rank model training as a warmup (Lialin et al., 2023), before optimizing in the low-rank subspace. There are two possible reasons: (1) the optimal weight matrices may not be low-rank, and (2) the reparameterization changes the gradient training dynamics.
@@
Notably, for pre-training, GaLore keeps low memory throughout the entire training, without requiring full-rank training warmup like ReLoRA. Thanks to GaLore’s memory efficiency, for the first time it is possible to train LLaMA 7B from scratch on a single GPU with 24GB memory (e.g., on NVIDIA RTX 4090), without any costly memory offloading techniques (Fig. 1).
@@ 2 Related Works
Low-Rank Adaptation
Hu et al. (2021) proposed Low-Rank Adaptation (LoRA) to fine-tune pre-trained models with low-rank adaptors. This method reduces the memory footprint by maintaining a low-rank weight adaptor for each layer. There are a few variants of LoRA proposed to enhance its performance (Renduchintala et al., 2023; Sheng et al., 2023; Xia et al., 2024), supporting multi-task learning (Wang et al., 2023), and further reducing the memory footprint (Dettmers et al., 2023). Lialin et al. (2023) proposed ReLoRA, a variant of LoRA designed for pre-training, but requires a full-rank training warmup to achieve comparable performance as the standard baseline.
Subspace Learning
Recent studies have demonstrated that the learning primarily occurs within a significantly low-dimensional parameter subspace (Larsen et al., 2022; Gur-Ari et al., 2018). These findings promote a special type of learning called subspace learning, where the model weights are optimized within a low-rank subspace. This notion has been widely used in different domains of machine learning, including meta-learning and continual learning (Lee & Choi, 2018; Chaudhry et al., 2020).
Projected Gradient Descent
@@
Memory-Efficient Optimization
There have been some works trying to reduce the memory cost of gradient statistics for adaptive optimization algorithms (Shazeer & Stern, ; Anil et al., ; Dettmers et al., 2021). Adafactor (Shazeer & Stern, ) achieves sub-linear memory cost by factorizing the second-order statistics by a row-column outer product. GaLore shares similarities with Adafactor in terms of utilizing low-rank factorization to reduce memory cost, but GaLore focuses on the low-rank structure of the gradients, while Adafactor focuses on the low-rank structure of the second-order statistics. GaLore can reduce the memory cost for both first-order and second-order statistics, and can be combined with Adafactor to achieve further memory reduction. Quantization is also widely used to reduce the memory cost of optimizer states (Dettmers et al., 2021; Li et al., 2023). Furthermore, Lv et al. (2023) proposed fused gradient computation to reduce the memory cost of storing weight gradients during training.
In contrast to the previous memory-efficient optimization methods, GaLore operates independently as the optimizers directly receive the low-rank gradients without knowing their full-rank counterparts.
@@ 3.1 Background
Low-rank updates.. For a linear layer $W\in\mathbb{R}^{m\times n}$, LoRA and its variants utilize the low-rank structure of the update matrix by introducing a low-rank adaptor $AB$:
@@ 3.2 Low-Rank Property of Weight Gradient
Surprisingly, while the weight matrices are not necessarily low-rank, the gradient indeed becomes low-rank during the training for certain gradient forms and associated network architectures:
@@
Let $m\leq n$ without loss of generality. The gradient update:
$G_{t}=A-BW_{t}C,\quad W_{t}=W_{t-1}+\eta G_{t-1}$ (6)
with constant $A$ and PSD matrices $B$ and $C$ and randomly initialized $W_{0}$ leads to low-rank gradient with high probability:
[…]
Here $\nu_{1}=\lambda_{\min}(C)$ is the smallest eigenvalues of $C$ and $\lambda_{1}\leq\ldots\leq\lambda_{n}$ are eigenvalues of $B$. Furthermore, if $\lambda_{2}>\lambda_{1}$ and $\nu_{1}>0$, then $G_{t}$ converges to rank-$1$ exponentially.
Note that in Lemma 3.1, we assume a parametric form (Eqn. 6) of the gradient. This is not a limiting assumption. It not only holds for simple linear network with objective $\varphi(W)=\|{\bm{y}}-W{\bm{x}}\|^{2}_{2}$, but also hold in more general nonlinear networks known as “reversible networks” (Tian et al., 2020), including deep ReLU networks:
Theorem 3.2 (Gradient Form of reversible models).
In a chained reversible neural network $\mathcal{N}({\bm{x}}):=\mathcal{N}_{L}(\mathcal{N}_{L-1}(\ldots\mathcal{N}_{1}({\bm{x}})))$ with $\ell_{2}$-objective $\varphi:=\frac{1}{2}\|{\bm{y}}-\mathcal{N}({\bm{x}})\|_{2}^{2}$, the weight matrix $W_{l}$ at layer $l$ has gradient $G_{l}$ of the following form for batchsize 1:
$G_{l}=\underbrace{J_{l}^{\top}{\bm{y}}{\bm{f}}_{l-1}^{\top}}_{A}-\underbrace{J^{\top}_{l}J_{l}}_{B}W_{l}\underbrace{{\bm{f}}_{l-1}{\bm{f}}_{l-1}^{\top}}_{C}$ (8)
where $J_{l}:=\mathrm{Jacobian}(\mathcal{N}_{L})\ldots\mathrm{Jacobian}(\mathcal{N}_{l+1})$ and ${\bm{f}}_{l}:=\mathcal{N}_{l}(\mathcal{N}_{l-1}\ldots\mathcal{N}_{1}({\bm{x}}))$.
Note that for softmax objective with small logits, we can also prove a similar structure of backpropagated gradient, and thus Theorem 3.2 can also apply.
Lemma 3.3 (Gradient structure of softmax loss).
With this lemma, it is clear that for a reversible network ${\bm{f}}:=\mathcal{N}({\bm{x}})=J_{l}({\bm{x}})W_{l}{\bm{f}}_{l-1}({\bm{x}})$, the gradient $G_{l}$ of $W_{l}$ has the following form:
which is consistent with the form $G_{l}=A-BW_{l}C$. For a detailed introduction to reversibility, please check the Appendix A.2.
@@ 3.3 Gradient Low-rank Projection (GaLore)
$W_{T}=W_{0}+\eta\sum_{t=0}^{T-1}\tilde{G}_{t},\quad\quad\tilde{G}_{t}=P_{t}\rho_{t}(P_{t}^{\top}G_{t}Q_{t})Q^{\top}_{t},$ (11)
@@
In the following, we show that GaLore converges under a similar (but more general) form of gradient update rule (Eqn. 6). This form corresponds to Eqn. 8 but with a larger batch size.
@@
Suppose the gradient has the following form (Eqn. 8 with batchsize $>1$):
$G=\sum_{i}A_{i}-\sum_{i}B_{i}WC_{i}$ (12)
where $B_{i}$ and $C_{i}$ are PSD matrices, $A_{i}$, $B_{i}$ and $C_{i}$ have $L_{A}$, $L_{B}$ and $L_{C}$ continuity with respect to $W$ and $\|W_{t}\|\leq D$. Let $R_{t}:=P_{t}^{\top}G_{t}Q_{t}$, $\hat{B}_{it}:=P_{t}^{\top}B_{i}(W_{t})P_{t}$, $\hat{C}_{it}:=Q_{t}^{\top}C_{i}(W_{t})Q_{t}$ and $\kappa_{t}:=\frac{1}{N}\sum_{i}\lambda_{\min}(\hat{B}_{it})\lambda_{\min}(\hat{C}_{it})$. If we choose constant $P_{t}=P$ and $Q_{t}=Q$, then GaLore with $\rho_{t}\equiv 1$ satisfies:
@@
Difference between GaLore and LoRA. While both GaLore and LoRA have “low-rank” in their names, they follow very different training trajectories. For example, when $r=\min(m,n)$, GaLore with $\rho_{t}\equiv 1$ follows the exact training trajectory of the original model, as $\tilde{G}_{t}=P_{t}P_{t}^{\top}G_{t}Q_{t}Q_{t}^{\top}=G_{t}$. On the other hand, when $BA$ reaches full rank (i.e., $B\in\mathbb{R}^{m\times m}$ and $A\in\mathbb{R}^{m\times n}$), optimizing $B$ and $A$ simultaneously follows very different training trajectory from the original model.

### v2

@@ Abstract
Training Large Language Models (LLMs) presents significant memory challenges, predominantly due to the growing size of weights and optimizer states. Common memory-reduction approaches, such as low-rank adaptation (LoRA), add a trainable low-rank matrix to the frozen pre-trained weight in each layer. However, such approaches typically underperform training with full-rank weights in both pre-training and fine-tuning stages since they limit the parameter search to a low-rank subspace and alter the training dynamics, and further, may require full-rank warm start. In this work, we propose Gradient Low-Rank Projection (GaLore), a training strategy that allows full-parameter learning but is more memory-efficient than common low-rank adaptation methods such as LoRA. Our approach reduces memory usage by up to 65.5% in optimizer states while maintaining both efficiency and performance for pre-training on LLaMA 1B and 7B architectures with C4 dataset with up to 19.7B tokens, and on fine-tuning RoBERTa on GLUE tasks. Our 8-bit GaLore further reduces optimizer memory by up to 82.5% and total training memory by 63.3%, compared to a BF16 baseline. Notably, we demonstrate, for the first time, the feasibility of pre-training a 7B model on consumer GPUs with 24GB memory (e.g., NVIDIA RTX 4090) without model parallel, checkpointing, or offloading strategies. Code is provided in the link.
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive performance across multiple disciplines, including conversational AI and language translation. However, pre-training and fine-tuning LLMs require not only a huge amount of computation but is also memory intensive. The memory requirements include not only billions of trainable parameters, but also their gradients and optimizer states (e.g., gradient momentum and variance in Adam) that can be larger than parameter storage themselves  (Raffel et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023). For example, pre-training a LLaMA 7B model from scratch with a single batch size requires at least 58 GB memory (14GB for trainable parameters, 42GB for Adam optimizer states and weight gradients, and 2GB for activations [^1] The calculation is based on LLaMA architecture, BF16 numerical format, and maximum sequence length of 2048.). This makes the training not feasible on consumer-level GPUs such as NVIDIA RTX 4090 with 24GB memory.
Figure 1: Estimated memory consumption of pre-training a LLaMA 7B model with a token batch size of 256 on a single device, without activation checkpointing and memory offloading [^2] In the figure, “no retaining grad” denotes the application of per-layer weight update to reduce memory consumption of storing weight gradient (Lv et al., 2023b).. Details refer to Section 5.5.
@@
Parameter-efficient fine-tuning (PEFT) techniques allow for the efficient adaptation of pre-trained language models (PLMs) to different downstream applications without the need to fine-tune all of the model’s parameters (Ding et al., 2022). Among them, the popular Low-Rank Adaptation (LoRA Hu et al. (2022)) reparameterizes weight matrix $W\in\mathbb{R}^{m\times n}$ into $W=W_{0}+BA$, where $W_{0}$ is a frozen full-rank matrix and $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$ are additive low-rank adaptors to be learned. Since the rank $r\ll\min(m,n)$, $A$ and $B$ contain fewer number of trainable parameters and thus smaller optimizer states. LoRA has been used extensively to reduce memory usage for fine-tuning in which $W_{0}$ is the frozen pre-trained weight. Its variant ReLoRA is also used in pre-training, by periodically updating $W_{0}$ using previously learned low-rank adaptors (Lialin et al., 2024).
However, many recent works demonstrate the limitation of such a low-rank reparameterization. For fine-tuning, LoRA is not shown to reach a comparable performance as full-rank fine-tuning (Xia et al., 2024). For pre-training from scratch, it is shown to require a full-rank model training as a warmup (Lialin et al., 2024), before optimizing in the low-rank subspace. There are two possible reasons: (1) the optimal weight matrices may not be low-rank, and (2) the reparameterization changes the gradient training dynamics.
@@
Notably, for pre-training, GaLore keeps low memory throughout the entire training, without requiring full-rank training warmup like ReLoRA. Thanks to GaLore’s memory efficiency, it is possible to train LLaMA 7B from scratch on a single GPU with 24GB memory (e.g., on NVIDIA RTX 4090), without any costly memory offloading techniques (Fig. 2).
@@ 2 Related Works
Low-rank adaptation.
Hu et al. (2022) proposed Low-Rank Adaptation (LoRA) to fine-tune pre-trained models with low-rank adaptors. This method reduces the memory footprint by maintaining a low-rank weight adaptor for each layer. There are a few variants of LoRA proposed to enhance its performance (Renduchintala et al., 2023; Sheng et al., 2023; Zhang et al., 2023; Xia et al., 2024), supporting multi-task learning (Wang et al., 2023b), and further reducing the memory footprint (Dettmers et al., 2024). Lialin et al. (2024) proposed ReLoRA, a variant of LoRA designed for pre-training, but requires a full-rank training warmup to achieve comparable performance as the standard baseline. Inspired by LoRA, Hao et al. (2024) also suggested that gradients can be compressed in a low-rank subspace, and they proposed to use random projections to compress the gradients. There have also been approaches that propose training networks with low-rank factorized weights from scratch (Kamalakara et al., 2022; Wang et al., 2023a; Zhao et al., 2023).
Subspace learning.
Recent studies have demonstrated that the learning primarily occurs within a significantly low-dimensional parameter subspace (Gur-Ari et al., 2018; Larsen et al., 2022). These findings promote a special type of learning called subspace learning, where the model weights are optimized within a low-rank subspace. This notion has been widely used in different domains of machine learning, including meta-learning and continual learning (Lee & Choi, 2018; Chaudhry et al., 2020).
Projected gradient descent.
@@
Low-rank gradient.
Gradient is naturally low-rank during training of neural networks, and this property have been studied in both theory and practice (Zhao et al., 2022; Cosson et al., 2023; Yang et al., 2023). It has been applied to reduce communication cost (Wang et al., 2018; Vogels et al., 2020), and memory footprint during training (Gooneratne et al., 2020; Huang et al., 2023; Modoranu et al., 2023).
Memory-efficient optimization.
There have been some works trying to reduce the memory cost of gradient statistics for adaptive optimization algorithms (Shazeer & Stern, 2018; Anil et al., 2019; Dettmers et al., 2022). Quantization is widely used to reduce the memory cost of optimizer states (Dettmers et al., 2022; Li et al., 2024). Recent works have also proposed to reduce weight gradient memory by fusing the backward operation with the optimizer update (Lv et al., 2023a, b).
@@ 3.1 Background
Low-rank updates. For a linear layer $W\in\mathbb{R}^{m\times n}$, LoRA and its variants utilize the low-rank structure of the update matrix by introducing a low-rank adaptor $AB$:
@@ 3.2 Low-Rank Property of Weight Gradient
Surprisingly, while the weight matrices are not necessarily low-rank, the gradient indeed becomes low-rank during the training for certain gradient forms and associated network architectures.
Reversible networks. Obviously, for a general loss function, its gradient can be arbitrary and is not necessarily low rank. Here we study the gradient structure for a general family of nonlinear networks known as “reversible networks” (Tian et al., 2020), which includes not only simple linear networks but also deep ReLU/polynomial networks:
Definition 3.1 (Reversiblity (Tian et al., 2020)).
A network $\mathcal{N}$ that maps input ${\bm{x}}$ to output ${\bm{y}}=\mathcal{N}({\bm{x}})$ is reversible, if there exists $L({\bm{x}};W)$ so that ${\bm{y}}=L({\bm{x}};W){\bm{x}}$, and the backpropagated gradient ${\bm{g}}_{\bm{x}}$ satisfies ${\bm{g}}_{\bm{x}}=L^{\top}({\bm{x}};W){\bm{g}}_{\bm{y}}$, where ${\bm{g}}_{\bm{y}}$ is the backpropagated gradient at the output ${\bm{y}}$. Here $L({\bm{x}};W)$ depends on the input ${\bm{x}}$ and weight $W$ in the network $\mathcal{N}$.
Please check Appendix B.1 for its properties. For reversible networks, the gradient takes a specific form.
Theorem 3.2 (Gradient Form of reversible models).
Consider a chained reversible neural network $\mathcal{N}({\bm{x}}):=\mathcal{N}_{L}(\mathcal{N}_{L-1}(\ldots\mathcal{N}_{1}({\bm{x}})))$ and define $J_{l}:=\mathrm{Jacobian}(\mathcal{N}_{L})\ldots\mathrm{Jacobian}(\mathcal{N}_{l+1})$ and ${\bm{f}}_{l}:=\mathcal{N}_{l}(\ldots\mathcal{N}_{1}({\bm{x}}))$. Then the weight matrix $W_{l}$ at layer $l$ has gradient $G_{l}$ in the following form for batch size 1:
(a) For $\ell_{2}$-objective $\varphi:=\frac{1}{2}\|{\bm{y}}-{\bm{f}}_{L}\|_{2}^{2}$:
$G_{l}=\left(J_{l}^{\top}{\bm{y}}-J^{\top}_{l}J_{l}W_{l}{\bm{f}}_{l-1}\right){\bm{f}}_{l-1}^{\top}$ (6)
(b) Left $P^{\perp}_{\bm{1}}:=I-\frac{1}{K}{\bm{1}}{\bm{1}}^{\top}$ be the zero-mean PSD projection matrix. For $K$-way logsoftmax loss $\varphi({\bm{y}};{\bm{f}}_{L}):=-\log\left(\frac{\exp({\bm{y}}^{\top}{\bm{f}}_{L})}{{\bm{1}}^{\top}\exp({\bm{f}}_{L})}\right)$ with small logits $\|P^{\perp}_{\bm{1}}{\bm{f}}_{L}\|_{\infty}\ll\sqrt{K}$:
$G_{l}=\left(J_{l}P^{\perp}_{\bm{1}}{\bm{y}}-\gamma K^{-1}J_{l}^{\top}P^{\perp}_{\bm{1}}J_{l}W_{l}{\bm{f}}_{l-1}\right){\bm{f}}_{l-1}^{\top}$ (7)
where $\gamma\approx 1$ and ${\bm{y}}$ is a data label with ${\bm{y}}^{\top}{\bm{1}}=1$.
From the theoretical analysis above, we can see that for batch size $N$, the gradient $G$ has certain structures: $G=\frac{1}{N}\sum_{i=1}^{N}(A_{i}-B_{i}WC_{i})$ for input-dependent matrix $A_{i}$, Positive Semi-definite (PSD) matrices $B_{i}$ and $C_{i}$. In the following, we prove that such a gradient will become low-rank during training in certain conditions:
@@
Suppose the gradient follows the parametric form:
$G_{t}=\frac{1}{N}\sum_{i=1}^{N}(A_{i}-B_{i}W_{t}C_{i})$ (8)
with constant $A_{i}$, PSD matrices $B_{i}$ and $C_{i}$ after $t\geq t_{0}$. We study vanilla SGD weight update: $W_{t}=W_{t-1}+\eta G_{t-1}$. Let $S:=\frac{1}{N}\sum_{i=1}^{N}C_{i}\otimes B_{i}$ and $\lambda_{1}<\lambda_{2}$ its two smallest distinct eigenvalues. Then the stable rank $\mathrm{sr}(G_{t})$ satisfies:
$\mathrm{sr}(G_{t})\leq\mathrm{sr}(G_{t_{0}}^{\parallel})\!+\!\left(\frac{1\!-\!\eta\lambda_{2}}{1\!-\!\eta\lambda_{1}}\right)^{2(t-t_{0})}\frac{\|G_{0}\!-\!G_{t_{0}}^{\parallel}\|_{F}^{2}}{\|G_{t_{0}}^{\parallel}\|_{2}^{2}}$ (9)
where $G_{t_{0}}^{\parallel}$ is the projection of $G_{t_{0}}$ onto the minimal eigenspace $\mathcal{V}_{1}$ of $S$ corresponding to $\lambda_{1}$.
In practice, the constant assumption can approximately hold for some time, in which the second term in Eq. 9 goes to zero exponentially and the stable rank of $G_{t}$ goes down, yielding low-rank gradient $G_{t}$. The final stable rank is determined by $\mathrm{sr}(G_{t_{0}}^{\parallel})$, which is estimated to be low-rank by the following:
Corollary 3.4 (Low-rank $G_{t}$).
If the gradient takes the parametric form $G_{t}=\frac{1}{N}\sum_{i=1}^{N}({\bm{a}}_{i}-B_{i}W_{t}{\bm{f}}_{i}){\bm{f}}_{i}^{\top}$ with all $B_{i}$ full-rank, and $N^{\prime}:=\operatorname{rank}(\{{\bm{f}}_{i}\})<n$, then $\mathrm{sr}(G_{t_{0}}^{\parallel})\leq n-N^{\prime}$ and thus $\mathrm{sr}(G_{t})\leq n/2$ for large $t$.
Remarks. The gradient form is justified by Theorem 3.2. Intuitively, when $N^{\prime}$ is small, $G_{t}$ is a summation of $N^{\prime}$ rank-1 update and is naturally low rank; on the other hand, when $N^{\prime}$ becomes larger and closer to $n$, then the training dynamics has smaller null space $\mathcal{V}_{1}$, which also makes $G_{t}$ low-rank. The full-rank assumption of $\{B_{i}\}$ is reasonable, e.g., in LLMs, the output dimensions of the networks (i.e., the vocabulary size) is often huge compared to matrix dimensions.
In general if the batch size $N$ is large, then it becomes a bit tricky to characterize the minimal eigenspace $\mathcal{V}_{1}$ of $S$. On the other hand, if $\mathcal{V}_{1}$ has nice structure, then $\mathrm{sr}(G_{t})$ can be bounded even further:
Corollary 3.5 (Low-rank $G_{t}$ with special structure of $\mathcal{V}_{1}$).
If $\mathcal{V}_{1}(S)$ is 1-dimensional with decomposable eigenvector ${\bm{v}}={\bm{y}}\otimes{\bm{z}}$, then $\mathrm{sr}(G_{t_{0}}^{\parallel})=1$ and thus $G_{t}$ becomes rank-1.
One rare failure case of Lemma 3.3 is when $G_{t_{0}}^{\parallel}$ is precisely zero, in which $\mathrm{sr}(G_{t_{0}}^{\parallel})$ becomes undefined. This happens to be true if $t_{0}=0$, i.e., $A_{i}$, $B_{i}$ and $C_{i}$ are constant throughout the entire training process. Fortunately, for practical training, this does not happen.
Transformers. For Transformers, we can also separately prove that the weight gradient of the lower layer (i.e., project-up) weight of feed forward network (FFN) becomes low rank over time, using the JoMA framework (Tian et al., 2024). Please check Appendix (Sec. B.3) for details.
@@ 3.3 Gradient Low-rank Projection (GaLore)
$W_{T}=W_{0}+\eta\sum_{t=0}^{T-1}\tilde{G}_{t},\quad\tilde{G}_{t}=P_{t}\rho_{t}(P_{t}^{\top}G_{t}Q_{t})Q^{\top}_{t}$ (10)
@@
In the following, we show that GaLore converges under a similar (but more general) form of gradient update rule (Eqn. 8). This form corresponds to Eqn. 6 but with a larger batch size.
@@
Suppose the gradient has the form of Eqn. 8 and $A_{i}$, $B_{i}$ and $C_{i}$ have $L_{A}$, $L_{B}$ and $L_{C}$ continuity with respect to $W$ and $\|W_{t}\|\leq D$. Let $R_{t}:=P_{t}^{\top}G_{t}Q_{t}$, $\hat{B}_{it}:=P_{t}^{\top}B_{i}(W_{t})P_{t}$, $\hat{C}_{it}:=Q_{t}^{\top}C_{i}(W_{t})Q_{t}$ and $\kappa_{t}:=\frac{1}{N}\sum_{i}\lambda_{\min}(\hat{B}_{it})\lambda_{\min}(\hat{C}_{it})$. If we choose constant $P_{t}=P$ and $Q_{t}=Q$, then GaLore with $\rho_{t}\equiv 1$ satisfies:
@@
Difference between GaLore and LoRA. While both GaLore and LoRA have “low-rank” in their names, they follow very different training trajectories. For example, when $r=\min(m,n)$, GaLore with $\rho_{t}\equiv 1$ follows the exact training trajectory of the original model, as $\tilde{G}_{t}=P_{t}P_{t}^{\top}G_{t}Q_{t}Q_{t}^{\top}=G_{t}$. On the other hand, when $BA$ reaches full rank (i.e., $B\in\mathbb{R}^{m\times m}$ and $A\in\mathbb{R}^{m\times n}$), optimizing $B$ and $A$ simultaneously follows a very different training trajectory compared to the original model.